OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 3x + y = 7 and passes through the point (6, −1). A. y = 1/3x − 3 B. y = 1/3x + 17 C. y = −3x − 3 D. y = −3x + 17

OpenStudy (anonymous):

Perpendicular lines have slopes that are negative reciprocals. What is the slope of 3x+y=7 ?

OpenStudy (anonymous):

the slope is got when you divide -x coefficient on y coefficient. so the slope is -3 and the slope of peripendicular line is 1/3; then y-(-1)=1/3(x-6) so A is true

OpenStudy (anonymous):

thanks!